OpenStudy (anonymous):

The moon forms a right triangle with the Earth and the sun during one of its phases, as shown below. A scientist measures the angle x and the distance y between the Earth and the moon. Please explain clearly how the scientist can use only these two measurements to calculate the distance between the Earth and the sun.

OpenStudy (anonymous):

here

OpenStudy (dumbcow):

you have to use a trig ratio: sin, cos, tan which one will use y and distance to sun but not the hypotenuse ?

OpenStudy (dumbcow):

sin = opp/hyp cos = adj/hyp tan = opp/adj -- only tan does not use hypotenuse, so to solve for distance (adjacent side) we need to use tangent --> tan(x) = y/d solve for d --> d = y/tan(x)

OpenStudy (anonymous):

@dumbcow a little bit clearer please

OpenStudy (dumbcow):

not sure how to be any clearer...im assuming you've learned trig ratios ?

OpenStudy (anonymous):

I'm super confused! I'm doing a practice sheet for my semester exam... I went over them but they didn't stick... :(

OpenStudy (dumbcow):

 an Acronym to use is SOH CAH TOA SOH- sin = opp/hyp CAH - cos = adj/hyp TOA - tan= opp/adj